The working substance for a Carnot cycle is 8 lb of air. The volume at the beginning of isothermal expansion is 9 cubic feet and the pressure is 300 psi. The ratio of expansion during the addition of heat is 2 and the temperature of the cold body is 90 degrees Fahrenheit . NOTE: Indicate the pV and TS planes Find: - QA(heat added) - QR(heat rejected) - MEP(cycle mean effective pressure) -the ratio of expansion during the isentropic process - the overall ratio of compression.
1
Expert's answer
2020-10-27T10:12:46-0400

n = number of moles

R = universal gas constant "= 8.3145 J\/mol K"

N = number of molecules

k = Boltzmann constant = "1.38066 \times 10-23 J/K = 8.617385 \times10-\frac{5 eV}{K}"

NA"= 8.3145 J\/mol K"

NA = Avogadro`s number "= 6.0221 \times 1023 /mol."